OpenStudy (anonymous):

Solve each equation over the interval [0, 2pi) tan^2x-3=0

OpenStudy (anonymous):

\[ \tan^2(x)= 3\\ \tan (x)= \pm \sqrt3\\ \] Can you finish it?

OpenStudy (anonymous):

Hint \[ \tan(\pi/3)=\sqrt 3 \]

OpenStudy (anonymous):

and also \[ \tan(-\pi/3)=-\sqrt 3 \] You should be able to finish it now

OpenStudy (anonymous):

We have to find the solutions that are in the interval \( [0 ,2 \pi)\)

OpenStudy (anonymous):

A moment of reflection reveals that \[ \frac \pi 3\\ \frac \pi 3+\pi=\frac {4 \pi}3\\ -\frac \pi 3+\pi=\frac {2 \pi}3\\ -\frac \pi 3+2\pi=\frac {5 \pi}3\\ \] I think that all the solutions.

OpenStudy (anonymous):

Thanks you! Sorry I keep losing.connection and its not letting me reply

OpenStudy (anonymous):

okay so tan^2x-3=0 You added 3 to both sides which turn out to be tan^2x= 3 then you square root both sides which make it cross the square root with tan^2 Which now equals tanx= square root of 3 Then you use the inverse trig function which would be x=inverse of tan square root of 3 + pi n Which equals x= 1.0479 + 2 pi And if you make a n and x table you get -2.0944 when n=1, 1.0479 when n= 0, 4.1888 when n=1 and 7.3304 when n=2 but since the interval is [0,2pi) the solutions are x=1.0479 and 4.1888 since 2pi= 6.2832 right??